A ball moving with a velocity v strike a wall moving towards the ball with a velocity u. An elastic impact lasts for t second.then the mean elastic force acting on the ball is

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Wasim
 
Relative speed of ball = (v+u)
speed after rebouncing = – (v+u)
 
F = m * dv/dt = m [v+ u + v + u]/t
 
F = 2m (v+ u) /t